solve the following system of equations using substitutions 2x-3y=-2 and y=-4x+24
Answer by mando271(42)

You are already given what y equals in the 2nd equation, just plug that into the 1st equation and solve for x.
2x-3(-4x+24)=-2
2x+12x-72=-2
14x=70
x=5
now plug this into the 2nd equation to solve for y.
y=-4(5)+24
y=-20+24
y=4
plug both numbers back into both orginal equations to verify.